Please help me...
If log2=x and log3=y,the value log60 is..

Answer by rothauserc(4718)

Using logarithm rule for multiplication
log60 = log6 +log10 = log2 +log3 +log10
log60 = log2 +log3 + log2 +log5
log60 = 2x +y + log5